# Spectro-Photometry and Radial Distribution of Multiple Stellar Populations in Globular Clusters from Gaia XP Spectra

### v1

@@ 3.2 Multiple populations in metal-poor GCs
In this subsection, we used the procedure that we have introduced for 47 Tucanae to derive photometry in the m38000100, m43000100 and m92250150 bands for stars in four GCs with different metallicities: NGC 6121, NGC 6752, NGC 3201, and NGC 6397. According to the 2010 edition of the Harris (1996) catalog, these clusters have iron abundances [Fe/H]=--1.16, --1.54, --1.59, and --2.02, respectively.
@@
The resulting $\Delta_{\rm m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ ChMs of RGB stars are plotted in the left and middle panels of Figure 5, whereas the right panels show the Na-O anticorrelation for NGC 6121 (from Marino et al., 2008) and the Al-Mg anticorrelations for NGC 6397, NGC 3201, and NGC 6752 (from APOGEE DR17). The stars for which both spectroscopy and photometry are available are represented with large dots and the probable 1P and 2P stars, selected from the Al-Mg or Na-O anticorrelations, are colored red and blue, respectively. The gray points show the remaining stars. Similarly with what is observed for 47,Tucanae, the bulk of 1P stars of all clusters are located around the origin of the ChM and have smaller average $\Delta_{\rm m3800,m4300,GRP}$ and larger $\Delta_{\rm m4300,m9225}$ absolute values than the 2P stars. These results indicate that the $\Delta_{\rm m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ ChM is an efficient tool to detect multiple populations in GCs with different light-element abundance. We derive similar conclusions by selecting the 1P and 2P stars from the [N/Fe] vs. [C/Fe] plane.
@@
The bulk of 1P and 2P stars populate distinct regions in the ChMs of NGC 6121, NGC 6752, and NGC 3201. On the contrary, the two stellar populations partially overlap each other in the ChM of NGC 6397. This fact is due to the low metallicity of this GC and the moderate light-element differences between its 1P and 2P stars.
@@
Figure 5: The left and middle panels show the $\Delta_{\rm m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ ChMs for RGB stars of NGC 6121, NGC 6752, NGC 3201, and NGC 6397. The right panels plot the anticorrelation between light elements. Specifically, we show [Al/Fe] vs. [Mg/Fe] for NGC 6752 and NGC 6397, and NGC 3201 and [Na/Fe] vs. [O/Fe] for NGC 6121. The abundances of aluminum and magnesium are taken from the from the 17th APOGEE data release, whereas sodium and oxygen abundances are derived by Marino et al. (2008). The probable 1P and 2P stars, for which both photometry and spectroscopy are available and that we selected on the basis of their light-element abundances, are colored red and blue, respectively, in the middle and right panels.
@@
Figure 6: This figure illustrates the procedure to derive the fractions of 1P and 2P stars in NGC 6121. The left panel reproduces the ChM of Figure 5, where we colored red the probable 1P stars. The aqua line represents the best fit straight line of the red stars and defines an angle $\theta=30^{o}$ with respect to the horizontal dashed line. The middle panel shows the rotated $\Delta_{2}$ vs. $\Delta_{1}$ diagram, whereas the histogram distribution of $\Delta_{2}$ is plotted in the right panel. The red curve is the best-fit Gaussian of the bona-fide 1P stars
@@ 4 Fraction and radial distribution of multiple populations
To estimate the fractions of 1P and 2P stars in each GC, we adopted the procedure illustrated in Figure 6 for NGC 6121, which has been already used in previous work on GCs (Milone et al., 2017). The aqua line represents a linear fit to the 1P stars. The $\Delta_{2}$ vs. $\Delta_{1}$ diagram plotted on the middle panel of Figure 6 has been generated by rotating counterclockwise the diagram in the left-hand panel by an angle of $\theta=30^{o}$, where $\theta$ is the angle between the green line and the dashed horizontal line. We derived the black histogram plotted in the right-hand panel, which represents the distribution of $\Delta_{2}$ for cluster stars. The distribution of bona-fide 1P stars is fitted with a Gaussian curve, represented by the red continuous line. The fraction of 1P stars is determined by calculating the ratio between the area under the Gaussian curve and the total number of RGB stars in the ChM. We found that the 1P of NGC 6121 comprises 30.0$\pm$3.1of the total numbers of studied RGB stars. Similarly, we derived the fractions of 1P stars along the RGB of 47 Tucanae, NGC 3201, and NGC 6752, which correspond to 48.8$\pm$2.4, 74.1$\pm$3.6, and 25.9$\pm$3.2%, respectively. We excluded NGC 6397 from the analysis because of the poor separation between 1P and 2P stars in the ChM of this cluster.
@@
To investigate the radial distribution of the multiple populations in each GC, we divided our sample into equally-populated groups of stars with different radial distances and estimated the fractions of 1P and 2P stars with the procedure discussed above. We adopted two radial bins for NGC 3201, three bins for NGC 6121 and NGC 6752, and four bins for 47 Tucanae.
@@
The results are illustrated in Figure 7. The fractions of 2P stars in both NGC 6121 and NGC 6752 are constant in the studied fields of view and are consistent with the values derived in the central regions from literature, thus indicating that the 1P and 2P stars are fully mixed in these clusters. In contrast, the fractions of 2P stars that we measured in the external regions of NGC 3201 and 47 Tucanae are much smaller than those observed in the cluster centers, thus indicating that their 2P stars are more centrally concentrated than the 1P.
@@
Figure 7: Fraction of 2P stars as a function of the radial distance from cluster center in 47 Tucanae, NGC 6121, NGC 6752, and NGC 3201. The results from this paper are represented with black diamonds, whereas literature determinations (Milone et al., 2013, 2014, 2017, 2019; Nardiello et al., 2015; Dondoglio et al., 2021, 2022; Marino et al., 2024) are plotted with light-red dots. The horizontal segments associated with each point mark the radial interval used to derive the corresponding fraction of 2P stars. The dotted, dashed, and dashed-dot lines indicate the core, half-light, and tidal radii, respectively, from the 2010 version of the Harris (1996) catalog.
@@ 5 Summary and discussion
We used Gaia DR3 low-resolution XP spectra to introduce new photometric bands that are sensitive to the different chemical compositions of multiple populations in GCs. To do this, we used simulated spectra of RGB stars in 47 Tucanae with the same metallicities but different light-element abundances as expected for 1P and 2P stars in GCs. We found that photometric diagrams constructed with magnitudes in the m38000200, m43000100, and m92250150 bands introduced in our work and in the GRP filter of Gaia DR3, are efficient tools to identify multiple populations among giant stars of 47 Tucanae. In particular, the $\Delta_{\rm m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ ChM allowed us to disentangle 1P and 2P stars along the RGB, red HB, and AGB and to detect star-to-star metallicity variations among 1P stars. The photometric identification of 1P and 2P stars is validated by the comparison of the ChM with the abundances of C, N, O, Mg, Al, and Na available from the literature and with the photometric diagrams constructed with the U,B,V,I magnitudes, which are traditionally used to identify multiple populations in GCs. Similarly, we clearly identified 1P and 2P stars along the RGBs of NGC 3201, NGC 6121, and NGC 6752. The sequences of 1P and 2P stars partially overlap each other in the ChM of NGC 6397, due to the small differences in the chemical composition of its multiple populations.
@@
The ChMs that we have introduced provide the opportunity to derive the fractions of 1P and 2P stars in the clusters outskirts, thus investigating the radial distribution of the multiple stellar populations near the tidal radius. The present day radial distribution of 1P and 2P stars in GCs can offer insights into their initial configurations and helps constrain the scenarios on the formation of multiple populations. There is widespread acceptance that in several GCs, such as 47 Tucanae, NGC 2808, NGC 5272, and NGC 5927, 2P stars are more centrally concentrated than 1P stars. In contrast, the radial distributions of multiple populations in some other clusters, including NGC 6752, NGC 6121, NGC 6366 and NGC 6838 remain consistent with one another (Dondoglio et al., 2021, 2022; Lee, 2019; Leitinger et al., 2023; Milone et al., 2012b, 2019; Nardiello et al., 2015; Simioni et al., 2016). These observations align with scenarios in which 2P stars originate in the cluster center and exhibit greater initial central concentration than the 1P stars. Consequently, clusters featuring persistently centrally concentrated 2P stars retain the memory of their initial distribution of multiple populations. In contrast, the 1P and 2P stars in other GCs become fully mixed over time due to dynamic evolution (e.g. Vesperini et al., 2013, 2018, 2021).
@@
This picture has been challenged by recent works that conclude that the 1P stars of certain GCs are more-centrally concentrated than the 2P, thus suggesting that they might actually form with more central concentrations. NGC 3201 stands out as a notable example illustrating this possibility (Leitinger et al., 2023).
@@
The ChMs that we derived from Gaia XP allow us to derive the fractions of 1P and 2P stars in the cluster outskirts, and explore a radial interval that is more extended than that analyzed in previous studies. Our findings, combined with literature results, confirm that the stellar populations of NGC 6121 and NGC 6752 share similar radial distributions. In contrast, the 2P stars of 47 Tucanae are more centrally concentrated than the 1P. The fraction of 2P stars declines from $\sim$80in the cluster core to $\sim$50at a radial distance of $\sim$3 half-light radii from the cluster center and remains nearly constant in the outermost cluster regions. These results are in agreement with the conclusions from previous studies (e.g. Milone et al., 2012a, 2019; Cordero et al., 2014; Lee, 2019, 2022; Dondoglio et al., 2021; Leitinger et al., 2023)
@@
Similarly to what we observe in 47 Tucanae, the 2P stars of NGC 3201 are centrally concentrated. Their fraction ranges from $\sim$55in the cluster core to $\sim$25between the half-light and the tidal radius. This result is in tension with the conclusions by Leitinger et al. (2023) that the 1P of NGC 3201 are more-centrally concentrated than 2P stars, and that the fraction of 2P stars is larger than $\sim$70in its outermost regions (Cadelano et al., 2024).
@@
Our results on the radial distribution of NGC 6121, NGC 6752, 47 Tucanae, and NGC 3201 are all consistent with the predictions of the scenarios for the GC formation where the 2P stars originate in the innermost cluster regions. We do not confirm the possibility suggested by Leitinger and collaborators and based on their conclusions on NGC 3201 that the 1P of some GCs can be more centrally concentrated than the 1P at the time of cluster formation. On the contrary, the radial distribution of its multiple stellar populations is comparable with that of other dinamically young star clusters, where the 2P is more centrally concentrated than the 1P.

### v2

@@ 3.2 Multiple populations in low-metallicity GCs
In this subsection, we apply the procedure that we have introduced for 47 Tucanae to derive photometry in the m38000100, m43000100 and m92250150 bands for stars in three GCs with different [Fe/H] values: NGC 6121, NGC 6752, and NGC 6397. According to the 2010 edition of the Harris (1996) catalog, these clusters have iron abundances [Fe/H]=--1.16, --1.54, and --2.02, respectively.
@@
The resulting $\Delta_{\rm m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ ChMs of RGB stars are plotted in the left and middle panels of Figure 7, whereas the right panels show the Na-O anticorrelation for NGC 6121 (from Marino et al., 2008) and the Al-Mg anticorrelations for NGC 6397, and NGC 6752 (from APOGEE DR17). The stars with both spectroscopic and photometric data are shown as large dots, with probable 1P and 2P stars, identified based on the Al-Mg or Na-O anticorrelations, colored red and blue, respectively. The gray points show the remaining stars. Similarly with what is observed for 47,Tucanae, the majority of 1P stars in all clusters are concentrated near the origin of the ChM and have smaller average $\Delta_{\rm m3800,m4300,GRP}$ and larger $\Delta_{\rm m4300,m9225}$ absolute values than the 2P stars. We found that the 1P and 2P stars of each cluster, identified from the ChM derived from the Gaia XP spectra populate distinct sequence of the $\Delta_{\rm{\it C}U,B,I}$ vs. $\Delta_{\rm B,I}$ ChM. Moreover, we verified that there is correspondence between the stellar populations selected from the $\Delta_{C\rm F336W,F438W,F814W}$ vs. $\Delta_{\rm F438W,F814W}$ and the $\Delta_{C\rm F275W,F336W,F438W}$ vs. $\Delta_{\rm F275W,F814W}$ ChMs.
These results indicate that the $\Delta_{\rm m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ ChM is an effective tool for detecting GC stars with different content of light-elements. We derive similar conclusions by selecting the 1P and 2P stars from the [N/Fe] vs. [C/Fe] plane.
@@
The bulk of 1P and 2P stars populate distinct regions in the ChMs of NGC 6121 and NGC 6752. On the contrary, the two stellar populations partially overlap each other in the ChM of NGC 6397. This fact is due to the low metallicity of this GC and the moderate light-element variations among its stars.
@@
Figure 7: $\Delta_{\rm m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ ChMs for RGB stars of NGC 6121, NGC 6752, and NGC 6397 (left and middle panels). The right panels plot the anticorrelation between light elements. Specifically, we show [Al/Fe] vs. [Mg/Fe] for NGC 6752 and NGC 6397, and [Na/Fe] vs. [O/Fe] for NGC 6121. The abundances of aluminum and magnesium are taken from the from the 17th APOGEE data release, whereas sodium and oxygen abundances are derived by Marino et al. (2008). In the middle and right panels, the probable 1P and 2P stars, for which both stellar magnitudes and spectroscopic data are accessible and identified based on their chemical compositions, are shown in red and blue, respectively.
Figure 8: $\Delta_{\rm m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ ChMs for RGB stars of NGC 3201 (top panels). The bottom panels show the Mg-Al and C-N anticorrelations from the 17th APOGEE data release. The 2P stars with large abundances of aluminum and nitrogen, are colored blue, whereas the remaining stars with available photometry and spectroscopy are represented with blue dots.
Figure 9: $I$ vs. $C_{\rm U,B,I}$ pseudo-CMD of NGC 3201 (left panel). The top-right panel is a zoom-in of the RGB region where the sequences of 1P and 2P stars are more evident for stars with radial distance from the cluster center larger than four arcmin. The bottom-right panels show the $\Delta_{\rm m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ (left) and the $\Delta_{C{\rm U,B,I}}$ vs. $\Delta_{\rm B,I}$ (right) ChMs. Probable 1P and 2P stars, selected from the ChM obtained from Gaia XP spectra, are colored red and blue, respectively.
Figure 10: $m_{\rm F814W}$ vs. $C_{\rm F336W,F438W,F814W}$ pseudo-CMD of NGC 3201 (left). Right panels compare the $\Delta_{C\rm F275W,F336W,F438W}$ vs. $\Delta_{\rm F275W,F814W}$ (top) and the $\Delta_{C\rm F336W,F438W,F814W}$ vs. $\Delta_{\rm F438W,F814W}$ (bottom) ChMs of RGB stars. The dashed gray line in the top-right ChM separates the bulk of 1P and 2P stars identified by Milone et al. (2017), which are colored red and blue, respectively, in all panels. The red triangles mark the 2P stars with low $\Delta_{C\rm F336W,F438W,F814W}$ values.
3.3 Multiple populations in NGC3201
In the following, we investigate NGC 3201 by using the same method adopted above for the other studied GCs. The $\Delta_{\rm m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ diagram for RGB stars is shown in the top panels of Figure 8 and reveals a group of stars with $\Delta_{\rm m3800,m4300,GRP}\gtrsim 0.2$ mag, alongside another group of stars with smaller $\Delta_{\rm m3800,m4300,GRP}$ values.
The bulk of probable 2P stars, selected from the Al-Mg and the C-N anticorrelations, are colored red and correspond to the group of stars with large values of $\Delta_{\rm m3800,m4300,GRP}$ in the ChM, whereas the remaining stars in the ChM are Al-poor and N-poor. Noticeably, at odds with what we observed in 47 Tucanae, NGC 6121, and NGC 6752, the group of ChM stars with low values of $\Delta_{\rm m3800,m4300,GRP}$ may exhibit intrinsic light-element variations, as suggested by the fact that the nitrogen abundance spans a range of $\sim$0.5 dex.
To compare the ChMs from Gaia-based and ground-based photometry, we take advantage from the photometric catalogs by Stetson et al. (2019). The photometry has been analyzed by following the same methods of our previous works (e.g.  Milone et al., 2018; Cordoni et al., 2020; Jang et al., 2022). In particular, we selected the cluster members from Gaia DR3 proper motions and parallaxes and corrected the photometry for the effects of differential reddening as in Milone et al. (2012a). The resulting $I$ vs. $C_{\rm U,B,I}$ is plotted in the left panel of Figure 9 for the NGC 3201 stars with radial distance larger than 1.5 arcmin. Moreover, in the bottom-right panels we compared the $\Delta_{\rm m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ and the $\Delta_{C{\rm U,B,I}}$ vs. $\Delta_{\rm B,I}$ ChMs (Jang et al., 2022). For completeness, in the top-right panel we show the $I$ vs. $C_{\rm U,B,I}$ diagrams for RGB stars with radial distances larger than 4.0 arcmin from the cluster center. The probable 2P stars, selected from the $\Delta_{\rm m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ ChM, are colored blue, while the remaining star are plotted in red. As expected, the selected stellar groups occupy distinct sequences in both the ChMs and the pseudo-CMD, confirming that these diagrams consistently identify the same stellar populations.
Finally, Figure 10, compares the $m_{\rm F814W}$ vs. $C_{\rm F336W,F438W,F814W}$ and $\Delta_{C\rm F336W,F438W,F814W}$ vs. $\Delta_{\rm F438W,F814W}$ diagrams, which are the HST-analogs of $I$ vs. $C_{\rm U,B,I}$ and $\Delta_{C\rm U,B,I}$ vs. $\Delta_{\rm B,I}$, with the $\Delta_{C\rm F275W,F336W,F438W}$ vs. $\Delta_{\rm F275W,F814W}$ diagram, the most commonly used ChM for studying multiple populations along the RGB. The dashed line superimposed on the latter ChM is derived by Milone et al. (2017) and separates the bulk of 1P and 2P stars. Clearly, all 1P stars (red circles) are located along the RGB sequence with low $C_{\rm F336W,F438W,F814W}$ values and cluster near the origin of the $\Delta_{C\rm F336W,F438W,F814W}$ vs. $\Delta_{\rm F438W,F814W}$ diagram. Moreover, all 2P stars (blue circles) have large values of $C_{\rm F336W,F438W,F814W}$ and $\Delta_{C\rm F336W,F438W,F814W}$. However, we notice a significant number of 2P stars (red triangles) that populated the RGB sequence with low $C_{\rm F336W,F438W,F814W}$ values and cluster near the origin of the $\Delta_{C\rm F336W,F438W,F814W}$ vs. $\Delta_{\rm F438W,F814W}$ diagram.
We conclude that the available photometric diagrams based on $U,B,I$ photometry allow us to disentangle 2P stars of NGC 3201 with large nitrogen variations (hereafter extreme 2P, or e2P stars), whereas the 2P stars that are mildly enhanced in nitrogen are mixed with the 1P. Due to the similar results obtained from $U,B,I$ Gaia-based photometry, the same conclusion can be extended to the $\Delta_{\rm m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ ChM.
@@ 3.3 Multiple populations in NGC3201
Figure 11: Method used to measure the fractions of 1P and 2P stars in NGC 6121. The left panel reproduces the ChM of Figure 7, where we colored red the probable 1P stars. The aqua line is the best-fit straight line through the red stars and forms an angle of $\theta=30^{o}$ with the gray dashed line. The middle panel displays the rotated $\Delta_{2}$ vs. $\Delta_{1}$ diagram, whereas the histogram distribution of $\Delta_{2}$ is plotted in the right panel. The red curve is the best-fit Gaussian of the 1P stars
@@ 4 Fraction and radial distribution of multiple populations
To estimate the fractions of 1P and 2P stars in each GC, we adopted the procedure illustrated in Figure 11 for NGC 6121, which has been already used in previous work on GCs (Milone et al., 2017). The aqua line represents a linear fit to the 1P stars. The $\Delta_{2}$ vs. $\Delta_{1}$ diagram plotted on the middle of Figure 11 was generated by rotating counterclockwise the ChM by $\theta=30^{o}$, where $\theta$ is traced from the green line to the dashed horizontal line. We obtained the histogram distribution of $\Delta_{2}$ for the cluster members and displayed it in the right-hand panel.
The distribution of bona-fide 1P stars is modeled using a Gaussian curve through least squares fitting. The fraction of 1P stars is estimated by taking the ratio of the area under the Gaussian curve (red line in Figure 11) to the total area of the histogram.
We found that the 1P of NGC 6121 comprises 30.0$\pm$3.1of the total numbers of studied RGB stars. Similarly, we derived the fractions of 1P stars along the RGB of 47 Tucanae, and NGC 6752, which correspond to 48.8$\pm$2.4, and 25.9$\pm$3.2%, respectively. In the case of NGC 3201, where it is not possible to disentangle 1P stars and 2P stars with moderate light-element variations, we derived the fraction of e2P stars, which corresponds to 25.9$\pm$3.6%. We excluded NGC 6397 from the analysis because 1P and 2P stars are partially mixed in its ChM.
@@
To analyze the radial distribution of multiple populations within each cluster, we segmented our sample into groups of stars with equal populations at varying radial distances and calculated the fractions of 2P stars (or e2P stars for NGC 3201) using the previously discussed method. We adopted two radial bins for NGC 3201, three bins for NGC 6121 and NGC 6752, and four bins for 47 Tucanae.
The results are illustrated in Figure 12, where we combined findings from this paper with those from previous studies, to analyze the radial behavior of the multiple populations.
In NGC 3201, we also take advantage from the $\Delta_{C{\rm U,B,I}}$ vs. $\Delta_{\rm B,I}$ ChM to fully cover the field of view within the cluster tidal radius. We used this diagram to compute the fractions of 2Pe stars in three different radial bins following the procedure illustrated in Figure 11 (see also Jang et al., 2022).
@@
The fractions of 2P stars in both NGC 6121 and NGC,6752 remain constant across the studied fields of view and are comparable to the values obtained from literature in the central regions, suggesting that their 1P and 2P stars are fully mixed. In contrast, the fractions of e2P and 2P stars that we measured in the external regions of NGC 3201 and 47 Tucanae, respectively, are much smaller than those observed in the cluster centers, thus indicating that their second populations are more concentrated toward the center compared to the 1P.
@@
Figure 12: Fraction of 2P stars as a function of the radial distance from cluster center in 47 Tucanae, NGC 6121, and NGC 6752, and fraction of e2P stars in NGC 3201. The results obtained from the photometry derived from Gaia XP spectra are represented with black diamonds, whereas literature determinations (Milone et al., 2013, 2014, 2017, 2019; Nardiello et al., 2015; Dondoglio et al., 2021, 2022; Marino et al., 2024) are plotted with light-red dots. The gray and light-red triangles show the fractions of e2P stars in NGC 3201 derived in this paper from $U,B,I$ and $F336W,F438W,F814W$ photometry, respectively. The horizontal segments associated with each point mark the radial interval used to derive the corresponding fraction of 2P stars. The dotted, dashed, and dashed-dot lines indicate the core, half-light, and tidal radii, respectively, from the 2010 version of the Harris (1996) catalog.
@@ 5 Summary and discussion
We used Gaia DR3 low-resolution XP spectra to introduce new photometric bands that are sensitive to the different chemical compositions of multiple populations in GCs. To do this, we used simulated spectra of RGB stars in 47 Tucanae with the same metallicities but different light-element abundances as expected for 1P and 2P stars in GCs. We found that diagrams based on photometry in the m38000200, m43000100, and m92250150 bands introduced in our work and in the GRP filter of Gaia DR3, are able to efficiently identify multiple populations among giant stars of 47 Tucanae.In particular, the $\Delta_{\rm m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ ChM enabled us to distinguish between 1P and 2P stars along the RGB, red HB, and AGB, as well as to identify metallicity variations among the 1P stars.
The photometric identification of 1P and 2P stars is validated by the comparison of the ChM with the abundances of C, N, O, Mg, Al, and Na available from the literature and with the photometric diagrams constructed with the U,B,V,I magnitudes, which are traditionally used to identify multiple populations in GCs. Similarly, we clearly identified 1P and 2P stars along the RGBs of NGC 3201, NGC 6121, and NGC 6752. The sequences of 1P and 2P stars in the ChM of NGC,6397 partially overlap, due to the small differences in the light-element abundances of its stars.
@@
The ChMs introduced in this paper enable the determination of 1P and 2P star fractions in the outskirts of clusters, allowing the investigation of the radial behavior of multiple stellar populations near the tidal radius. The current distribution of 1P and 2P stars in GCs provides insights into their initial configurations and helps constrain models of GC formation. There is widespread acceptance that in several GCs, such as 47 Tucanae, NGC 2808, NGC 5272, and NGC 5927, 2P stars exhibit a higher central concentration compared to 1P stars. In contrast, the radial distributions of multiple populations in some other clusters, including NGC 6752, NGC 6121, NGC 6366 and NGC 6838 remain consistent with one another (Dondoglio et al., 2021, 2022; Lee, 2019; Leitinger et al., 2023; Milone et al., 2012b, 2019; Nardiello et al., 2015; Simioni et al., 2016). These observations align with scenarios in which 2P stars originate in the cluster center and exhibit greater initial central concentration than the 1P stars. Consequently, clusters featuring persistently centrally concentrated 2P stars retain the imprint of the original stellar distribution. In contrast, the 1P and 2P stars in other GCs become thoroughly mixed over time as a result of dynamical evolution (e.g. Vesperini et al., 2013, 2018, 2021).
@@
This view has been challenged by recent studies, which suggest that in certain GCs, 2P stars are less centrally clustered than 1P stars. This conclusion implies that 1P stars might actually form with more central concentrations. NGC 3201 stands out as a notable example illustrating this possibility (Leitinger et al., 2023).
@@
The ChMs that we derived from Gaia XP allow us to derive the fractions of 1P and 2P stars in the cluster outskirts, and explore a radial interval that is more extended than that analyzed in previous studies. Our findings, combined with literature results, confirm that the stellar populations of NGC 6121 and NGC 6752 share similar radial distributions. In contrast, 47 Tucanae presents a more centrally concentrated distribution of 2P stars compared to the 1P. The fraction of 2P stars declines from $\sim$80in the cluster core to $\sim$50at a radial distance of $\sim$3 half-light radii and remains nearly constant in the outermost cluster regions. These results are consistent with the findings from previous studies (e.g. Milone et al., 2012a, 2019; Cordero et al., 2014; Lee, 2019, 2022; Dondoglio et al., 2021; Leitinger et al., 2023).
@@
Similarly to what we observe in 47 Tucanae, the e2P stars of NGC 3201 are centrally concentrated. Their fraction ranges from $\sim$55in the cluster core to $\sim$25between the half-light and the tidal radius. This result is in tension with the conclusions by Leitinger et al. (2023) that the 1P of NGC 3201 are more-centrally concentrated than 2P stars. Our findings are in qualitative agreement with those by Cadelano et al. (2024) for radial distances smaller than $\sim 4$ arcmin from the cluster center. However, we do not confirm the possibility of a bimodal radial distribution suggested by these authors with a fraction of e2P stars larger than $\sim$70in its outermost regions.
@@
Our results on the radial distribution of NGC 6121, NGC 6752, 47 Tucanae, and NGC 3201 are all consistent with the predictions of the scenarios for the GC formation where the 2P stars originate in the innermost cluster regions. We do not confirm the possibility suggested by Leitinger and collaborators and based on their conclusions on NGC 3201 that the 1P of some GCs can be more centrally clustered than the 1P at the epoch of cluster birth. On the contrary, the radial distribution of its multiple stellar populations is comparable with that of other dinamically young star clusters, where the 2P has a higher degree of central concentration than the 1P.